Give an example of a research question that would be suitable for computing a:

a) Correlation coefficient

b) Regression line
Answer #1

Answer:

Regression and correlation cofficient strategies are utilized to think about the connections between factors. Relapse is utilized to foresee the estimation of one variable dependent on the estimation of an alternate variable. Connection is a proportion of the quality of a connection between factors. The factors are information which are estimated and additionally included in an examination. On account of the precedents utilized here, the information were gotten by checking the breathing rate of goldfish in a lab test.

The word relationship is utilized in regular daily existence to mean some type of affiliation. We may state that we have seen a connection between's foggy days and assaults of wheeziness. Be that as it may, in measurable terms we use connection to signify relationship between two quantitative factors. We likewise accept that the affiliation is straight, that one variable increments or diminishes a fixed sum for a unit increment or abatement in the other. The other strategy that is frequently utilized in these conditions is relapse, which includes assessing the best straight line to abridge the affiliation.

The regression line

As noted over, a straight line plotted on the Cartesian organize framework can have the condition y = mx + b. We recollect that m is the incline and b is the y-block.

A regression line will have a general structure

y = a + bx + e

where:

an is the y-block

b is the incline of the line

e is a blunder term

Practically speaking, under conventional conditions, we don't have the foggiest idea about the estimation of the blunder term so we utilize the accompanying type of the condition

y = a + bx

albeit elective structures, (for example, y + hatchet + b) will likewise yield similar outcomes.

Regression line and correlation cofficient  relies upon a lot of estimations. These are finished by taking the separation that a point is from the hypothetical relapse line and squaring it. By including these squares you acquire the entirety of squares. Entirety of squares data can be controlled by computing essential measurements on the information of the reliant and free factors. Allude to the area on counts for relapse and relationship.